OpenStudy (anonymous):

Solve: 2x + 36 + 40 = 180

OpenStudy (anonymous):

36+40=74 Subtract both sides by this, then divide both sides by two. Got your answer now? :-)

OpenStudy (anonymous):

180 - 40 = 140 140 - 36 = 104 104/2 = 52 x = 52